Edwards Construction currently has debt outstanding with a market value of $93,000 and a cost of 8 percent. The company has EBIT of $7,440 that is expected to continue in perpetuity. Assume there are no taxes.

  

a-1.

What is the value of the company's equity? (Do not round intermediate calculations. Leave no cell blank - be certain to enter "0" wherever required.)

a-2. What is the debt-to-value ratio? (Do not round intermediate calculations and round your answer to the nearest whole number, e.g., 32.)
b. What are the equity value and debt-to-value ratio if the company's growth rate is 3 percent? (Do not round intermediate calculations and round your "Debt-to-value" answer to 3 decimal places, e.g., 32.161.)
c. What are the equity value and debt-to-value ratio if the company's growth rate is 7 percent? (Do not round intermediate calculations and round your "Debt-to-value" answer to 3 decimal places, e.g., 32.161.)
Answer #1

a-1). Interest = Debt * Interest Rate = $93,000 * 0.08 = $7,440

Cash Flow for shareholders = EBIT - Interest = $7,440 - $7,440 = $0

Value of Equity = $0

a-2). Debt-to-value ratio = Debt / Value of firm = $93,000 / $93,000 = 1.0

b). Earnings next year = Current Earnings * (1 + g) = $7,440 * (1 + 0.03) = $7,663.20

Since, there is no risk,the required return for shareholders is the same as the required return on the company’s debt. The payments to stockholders will increase at the growth rate of 3% (a growing perpetuity), so the value of these payments today is

Value of Firm = Earnings next year / (r - g) = $7,663.20 / (0.08 - 0.03) = $153,264

Value of Debt = $93,000

Value of Equity = Value of firm - Value of Debt = $153,264 - $93,000 = $60,264

Debt-to-value ratio = Debt / Value of firm = $93,000 / $153,264 = 0.6068, or 60.680%

c). Earnings next year = Current Earnings * (1 + g) = $7,440 * (1 + 0.07) = $7,960.80

Value of Firm = Earnings next year / (r - g) = $7,960.80 / (0.08 - 0.07) = $796,080

Value of Debt = $93,000

Value of Equity = Value of firm - Value of Debt = $796,080 - $93,000 = $703,080

Debt-to-value ratio = Debt / Value of firm = $93,000 / $796,080 = 0.11682, or 11.682%